Our sun has a wavelength of maximum intensity of emission approximately 510 nm. Use Wien’s law to estimate its surface temperature in Kelvin, to two significant figures. 


1
Expert's answer
2021-04-16T07:27:03-0400

According to Wien’s law:


"\lambda_\text{max}=\frac bT,\\\space\\
T=\frac{b}{\lambda_\text{max}}=\frac{2.898·10^{-3}}{510·10^{-9}}=5700\text{ K}."